OpenStudy (anonymous):

How many grams are in an ounce? This summer, the fire department is buying lighting bugs. They are paying $13 dollars an ounce. I read somewhere that a lightening bug is .15 grams but I keep getting confused when I try to convert it. I'm trying to figure out how many lightening bugs I would have to catch to make one ounce.

jimthompson5910 (jim_thompson5910):

1 gram = 0.035274 ounces (according to google's conversion) 0.15*1 gram = 0.15*0.035274 ounces ... multiply both sides by 0.15 0.15 gram = 0.0052911ounces ... multiply

jimthompson5910 (jim_thompson5910):

1 lightening bug = 0.15 grams 1 lightening bug = 0.0052911 ounces ... substitution 1 lightening bug/0.0052911 = 0.0052911 ounces/0.0052911 .. divide both sides by 0.0052911 1/0.0052911 = 0.0052911/0.0052911 188.996616960557 lightening bugs = 1 ounce 1 ounce = 188.996616960557 lightening bugs

jimthompson5910 (jim_thompson5910):

so you need roughly 189 lightening bugs (you round up to make sure you don't go under 1 ounce)

OpenStudy (anonymous):

oh wow, thanks for the details!

jimthompson5910 (jim_thompson5910):

yw